Find all solutions of the recurrence relation a_n=5a_(n-1)-6a_(n-2)+2^n+3n.
1
Expert's answer
2020-06-09T18:33:43-0400

Let f(z) generating function. f(z) = "\sum" an*zn and if n < 0 then an = 0

an = 5*an-1 -n6*an-2 + 2n + 3*n multiply by zn and add up

"\sum" 2n*zn = 1/(1-2z) , "\sum" 3*n*zn = 3z/(1-z)2 =>

f(z) = 5*f(z)*z -6*f(z)*z2 + 1/(1 - 2z) + 3z/(1-z)2 + b + c*z =>

a0 = f(0) = 1 + b , a1 = f'(0) =5*f(0)+ 2 + 3 = 5*a0 + 5 + c =>

b = a0 - 1 , c = a1 - 5a0 - 5 =>

f(z) = (1+3z)/((1-2z)(1-3z)(1-z)2) + (a0 - 1 + (a1 - 5a0 - 5)*z)/((1-2z)(1-3z)) =>

an = (a0 - 1)*(3n+1 - 2n+1) + (a1 - 5a0 - 5)*(3n - 2n) + (15 - 5*2n+3 + 3n+3 + 4n)/2